# High-dimensional analysis of ridge regression for non-identically distributed data with a variance profile

### v3

@@ A.3 Proof of the main results
Since Assumption 1.3 hold true for the variance profile $\widetilde{S}_{p}$ and $\Upsilon_{n}$, it follows that $(\widetilde{S}_{p})_{p\geq 1}$ and $(\widehat{\Sigma}_{n})_{n\geq 1}$ satisfy (3.3). Moreover the family of identity matrices, namely $(I_{p})_{p\geq 1}$, also satisfies (3.3) thus, we deduce from Theorem 3.1 that
$\frac{1}{p}\operatorname{Tr}[Q_{p}(z)\widetilde{S}_{p}]\sim\frac{1}{p}\operatorname{Tr}[T_{p}(z)\widetilde{S}_{p}]$ , $\frac{1}{p}\operatorname{Tr}[Q_{p}^{\prime}(z)\widetilde{S}_{p}]\sim\frac{1}{p}\operatorname{Tr}[T^{\prime}(z)\widetilde{S}_{p}]$ (A.3)
$\frac{1}{p}\operatorname{Tr}[Q_{p}^{\prime}(z)\widehat{\Sigma}_{n}]\sim\frac{1}{p}\operatorname{Tr}[T_{p}^{\prime}(z)\widehat{\Sigma}_{n}]$ , $\frac{1}{p}\operatorname{Tr}[Q_{p}^{\prime}(z)]\sim\frac{1}{p}\operatorname{Tr}[T^{\prime}(z)].$ (A.4)
Hence Equation (3.6) directly follows from (A.3) and (3.5) which concludes the proof of Theorem 3.2. ∎
@@
Under the additional Assumptions 1.4 and 1.5, [AEK17b][Theorem 2.9] asserts that there exists $\tau>0$ such that $\pi((0,\tau])=0$ and if $p<n$ then $\pi_{\ast}=0$. Since $T_{p}(-\lambda)=\int_{\mathbb{R}^{+}}\frac{\mu(dw)}{w+\lambda}$ for $\lambda>0$, one get that $\int_{0}^{\tau}\frac{\frac{1}{p}\mathrm{Tr}(\mu(dw))}{w+\lambda}=\int_{0}^{\tau}\frac{\nu_{p}(dw)}{w+\lambda}=0$.
@@
Moreover, $|\frac{1}{w+\lambda}|\leq\frac{1}{\tau}$ for $\lambda>0$ and $w\geq\tau$, thus by the dominated convergence theorem $T_{p}(-\lambda)$ admits a limit when $\lambda$ tends to 00, we denote $T_{p}(0^{-})=\int_{\tau}^{+\infty}\frac{\mu(dw)}{w}$ this limit. A similar proof allows us to state that $T_{p}^{\prime}(-\lambda)$ admits a limits when $\lambda$ tends to 00, we denote $T^{\prime}(0^{-})=\int_{\tau}^{+\infty}\frac{\mu(dw)}{w^{2}}$ this limit. ∎
@@
As stated in Section 2.3, in the case of a quasi doubly stochastic variance profile, $T_{p}(z)$ is a scalar matrix. Especially, $T_{p}(z)=m_{p}(z)I_{p}$ where $m_{p}(z)$ is the Stieltjes transform of the Marchenko-Pastur distribution. Moreover since $\frac{1}{p}\mathrm{Tr}[T_{p}(z)]=\int_{\mathbb{R}^{+}}\frac{\nu_{p}(dw)}{w+\lambda}$, we get $m_{p}(z)=\int_{\mathbb{R}^{+}}\frac{\nu_{p}(dw)}{w+\lambda},$ with $\nu_{p}(dw)=\pi_{\ast}\delta_{0}(dw)+\pi(dw)\mathbf{1}_{w>0}dw$. We have seen in the previous proof that there exists $\tau>0$ such that $\nu_{p}((0,\tau])=0$ and if $p<n$ then $\pi_{\ast}=0$. This implies that 00 belongs to the domain of definition of $m_{p}(.)$. Hence $\lim\limits_{\lambda\rightarrow 0}T_{p}(-\lambda)=\lim\limits_{\lambda\rightarrow 0}m_{p}(-\lambda)=m_{p}(0)$ and consequently $T_{p}(0^{-})=m_{p}(0)$.
We prove symmetrically that if $p>n$ then 00 belongs to the domain of definition of $\tilde{m}_{n}(.)$. Moreover, we have seen in Section 3.1 that in the case of a quasi doubly stochastic variance profile $\kappa(z)=\frac{1}{\tilde{m}_{n}(z)}I_{n}$. Note that $\tilde{m}_{n}(0)=\int_{\tilde{\tau}}^{+\infty}\frac{\tilde{\nu}_{n}(dw)}{w}$ is positive. Hence we conclude that $\kappa(0^{+})=\frac{1}{\tilde{m}_{n}(0)}I_{n}$. We prove $\kappa^{\prime}(0^{+})=\frac{\tilde{m}_{n}^{\prime}(0)}{\tilde{m}_{n}^{2}(0)}I_{n}$ the same way. ∎
@@
where $\mu=(\mu_{ij})$ is a $p\times p$ matrix valued measure such that $\mu(\mathbb{R}^{+})=I_{p}$. Since $w\mapsto\frac{1}{w/\lambda+1}$ is a measurable function for $\lambda>0$,$\lim\limits_{\lambda\rightarrow+\infty}\frac{1}{w/\lambda+1}=1$ for $w\in\mathbb{R}^{+}$ and $|\frac{1}{w/\lambda+1}|\leq 1$, we deduce from the dominated convergence theorem that $\int_{\mathbb{R}^{+}}\frac{\mu(dw)}{w/\lambda+1}\xrightarrow[\lambda\rightarrow+\infty]{}I_{p}.$ Which finally gives us
@@
[…]
@@
Moreover, according to Lemma 3.2 $\kappa(\lambda)$ and $\kappa^{\prime}(\lambda)$ admit limits when $\lambda$ tends to 00. Thus we finish this proof by getting considering the definitions of $\kappa(0^{+})$ and $\kappa^{\prime}(0^{+})$ from Lemma 3.2.
@@
Let us prove Corollary 3.3. The function $g(\lambda)=\hat{r}_{\lambda}^{test}(X_{n})$ is differentiable for $\lambda>0$, with $g^{\prime}(\lambda)=2\Big{(}\frac{\lambda\alpha^{2}}{p}-\frac{\sigma^{2}}{n}\Big{)}\mathrm{Tr}[\Sigma_{n}\widehat{\Sigma}_{n}Q_{p}(\lambda)^{3}].$ Since the symmetric and positive semi-definite matrices $\widehat{\Sigma}_{n}$ and $Q_{p}(\lambda)^{3}$ are commuting, then $\Sigma_{n}\hat{\Sigma}_{n}Q_{p}(\lambda)^{3}$ is also symmetric positive semi-definite.
Thus $\mathrm{Tr}[\Sigma_{n}\widehat{\Sigma}_{n}Q_{p}(\lambda)^{3}]\geq 0$ which means that the sign of $g^{\prime}(\lambda)$ only depends on $\Big{(}\frac{\lambda\alpha^{2}}{p}-\frac{\sigma^{2}}{n}\Big{)}$. Indeed, if $\lambda\leq\lambda_{\ast}$ then $g^{\prime}(\lambda)\leq 0$, and if $\lambda\geq\lambda_{\ast}$ then $g^{\prime}(\lambda)\geq 0$. This proves that Corollary 3.3 holds true, and this concludes the proof of Corollary 3.3. ∎
@@ A.4 Random matrices with a variance profile in free probability
and it is sometimes defined as $-\mathcal{G}_{A}$. The operator-valued $\mathcal{R}$-transform $\mathcal{R}_{A}$ of a Hermitian matrix $A$ is the unique analytic map satisfying,
@@
for all $\mathcal{Z}$ in $\mathbb{D}_{p}(\mathbb{C})^{-}$ whose diagonal entries have imaginary parts large enough in absolute value. An efficient way to produce a deterministic equivalent for a random matrix $A$ is then to find a simple approximation $\mathcal{R}_{A}^{\square}$ of $\mathcal{R}_{A}$. In this manner, one can then approximate the operator-valued Stieltjes transform of $A$ by the solution of the fixed point equation (A.9) where $\mathcal{R}_{A}$ is replaced by $\mathcal{R}_{A}^{\square}$. This method also allows us to compute the spectrum of perturbations of $A$ by independent matrices, see for instance [BM20].
For a generalized Wigner matrix $W_{p}$ of size $p\times p$ with a symmetric variance profile $\Gamma_{p}$, Shlyakhtenko proves [Shl96] that a good approximation of $\mathcal{R}_{W_{p}}$ is the deterministic linear map
$\mathcal{R}_{W_{p}}^{\square}(\mathcal{Z})=\mathrm{deg}\left(\frac{1}{p}\Gamma_{p}\mathcal{Z}\right),\quad\mbox{for all}\quad\mathcal{Z}\in\mathbb{D}_{p}(\mathbb{C})^{-},$
where for a matrix $A$, we denote by $\mathrm{deg}(A)$ the diagonal matrix, whose $k$-diagonal element is the sum of the entries of the $k$-row of $A$.
To the best of our knowledge, apart from generalized Wigner matrices, there does not exist any other class of random matrices for which a simple approximation of the diagonal-valued $\mathcal{R}$-transform is known yet. Nevertheless, we ca now remark that, for $\mathcal{Z}=zI_{p}$ with $z\in\mathbb{C}\setminus{\mathbb{R}}^{+}$, the diagonal matrix $T_{p}(z)$ solution of the Dyson equation (2.3) can be written as satisfying the fixed-point equation
$T_{p}(z)=\left(-zI_{p}+\mathcal{R}_{\Sigma_{n}}^{\square}(-T_{p}(z))\right)^{-1},$
where $\mathcal{R}_{\Sigma_{n}}^{\square}$ is the non-linear map
$\mathcal{R}_{\Sigma_{n}}^{\square}(\mathcal{Z})=\mathrm{deg}\left(\frac{1}{n}\Gamma_{n}^{\top}\left[I_{n}-\mathrm{deg}\left(\frac{1}{n}\Gamma_{n}\mathcal{Z}\right)\right]^{-1}\right),\quad\mbox{for all}\quad\mathcal{Z}\in\mathbb{D}_{p}(\mathbb{C})^{-}.$
This suggests that $\mathcal{R}_{\Sigma_{n}}^{\square}$ is indeed a simple approximation of the operator-valued $\mathcal{R}$-transform of $\Sigma_{n}$.

### v4

@@ A.3 Proof of the main results
Since Assumption 1.3 hold true for the variance profile $\widetilde{S}_{p}$ and $\Upsilon_{n}$, it follows that $(\widetilde{S}_{p})_{p\geq 1}$ and $(\widehat{\Sigma}_{n})_{n\geq 1}$ satisfy Equation (3.3). Moreover the family of identity matrices, namely $(I_{p})_{p\geq 1}$, also satisfies Equation (3.3) thus, we deduce from Theorem 3.1 that
[…]
$\frac{1}{p}\operatorname{Tr}[Q_{p}^{\prime}(z)\widehat{\Sigma}_{n}]\sim\frac{1}{p}\operatorname{Tr}[T_{p}^{\prime}(z)\widehat{\Sigma}_{n}]\mbox{ , }\frac{1}{p}\operatorname{Tr}[Q_{p}^{\prime}(z)]\sim\frac{1}{p}\operatorname{Tr}[T^{\prime}(z)].$ (A.4)
Hence Equation (3.6) directly follows from Equations (A.3) A.4 and (3.5). ∎
@@
Under the additional Assumptions 3.1 and 3.2, [AEK17b][Theorem 2.9] asserts that there exists $\tau>0$ such that $\pi((0,\tau])=0$ and if $p<n$ then $\pi_{\ast}=0$. Since $T_{p}(-\lambda)=\int_{\mathbb{R}^{+}}\frac{\mu(dw)}{w+\lambda}$ for $\lambda>0$, one get that $\int_{0}^{\tau}\frac{\frac{1}{p}\mathrm{Tr}(\mu(dw))}{w+\lambda}=\int_{0}^{\tau}\frac{\nu_{p}(dw)}{w+\lambda}=0$.
@@
Moreover, $|\frac{1}{w+\lambda}|\leq\frac{1}{\tau}$ for $\lambda>0$ and $w\geq\tau$, thus by the dominated convergence theorem $T_{p}(-\lambda)$ admits a limit when $\lambda$ tends to 0, we denote $T_{p}(0^{-})=\int_{\tau}^{+\infty}\frac{\mu(dw)}{w}$ this limit. A similar proof allows us to state that $T_{p}^{\prime}(-\lambda)$ admits a limits when $\lambda$ tends to 0, we denote $T^{\prime}(0^{-})=\int_{\tau}^{+\infty}\frac{\mu(dw)}{w^{2}}$ this limit. ∎
@@
As stated in Section 2.3, in the case of a quasi doubly stochastic variance profile, $T_{p}(z)$ is a scalar matrix. Especially, $T_{p}(z)=m_{p}(z)I_{p}$ where $m_{p}(z)$ is the Stieltjes transform of the Marchenko-Pastur distribution. Moreover, we know from Proposition 2.1 that $\frac{1}{p}\mathrm{Tr}[T_{p}(z)]$ is a Stieltjes Transform associated to a measure $\nu_{p}$. Therefore, $\frac{1}{p}\mathrm{Tr}[T_{p}(z)]=\int_{\mathbb{R}^{+}}\frac{\nu_{p}(dw)}{w+\lambda}$ and $m_{p}(z)=\int_{\mathbb{R}^{+}}\frac{\nu_{p}(dw)}{w+\lambda}.$ Additionnaly, one has from [AEK17b][Theorem 2.1] that $\nu_{p}(dw)=\pi_{\ast}\delta_{0}(dw)+\pi(dw)\mathbf{1}_{w>0}dw$. We have seen in the previous proof that there exists $\tau>0$ such that $\nu_{p}((0,\tau])=0$ and if $p<n$ then $\pi_{\ast}=0$. This implies that 0 belongs to the domain of definition of $m_{p}(.)$. Hence $\lim\limits_{\lambda\rightarrow 0}T_{p}(-\lambda)=\lim\limits_{\lambda\rightarrow 0}m_{p}(-\lambda)=m_{p}(0)$ and consequently $T_{p}(0^{-})=m_{p}(0)$.
We prove symmetrically that if $p>n$ then 0 belongs to the domain of definition of $\tilde{m}_{n}(.)$. Moreover, we have seen in Section 3.1 that in the case of a quasi doubly stochastic variance profile $\kappa(z)=\frac{1}{\tilde{m}_{n}(z)}I_{n}$. Note that $\tilde{m}_{n}(0)=\int_{\tilde{\tau}}^{+\infty}\frac{\tilde{\nu}_{n}(dw)}{w}$ is positive. Hence we conclude that $\kappa(0^{+})=\frac{1}{\tilde{m}_{n}(0)}I_{n}$. We prove $\kappa^{\prime}(0^{+})=\frac{\tilde{m}_{n}^{\prime}(0)}{\tilde{m}_{n}^{2}(0)}I_{n}$ the same way. ∎
@@
where $\mu=(\mu_{ij})$ is a $p\times p$ matrix valued measure such that $\mu(\mathbb{R}^{+})=I_{p}$. Since $w\mapsto\frac{1}{w/\lambda+1}$ is a measurable function for $\lambda>0$$\lim\limits_{\lambda\rightarrow+\infty}\frac{1}{w/\lambda+1}=1$ for $w\in\mathbb{R}^{+}$ and $|\frac{1}{w/\lambda+1}|\leq 1$, we deduce from the dominated convergence theorem that $\int_{\mathbb{R}^{+}}\frac{\mu(dw)}{w/\lambda+1}\xrightarrow[\lambda\rightarrow+\infty]{}I_{p}.$ Which finally gives us
@@
$\lambda T_{p}^{\prime}(-\lambda)\xrightarrow[\lambda\rightarrow+\infty]{}0_{p}\quad\mathrm{and}\quad\lambda^{2}T_{p}^{\prime}(-\lambda)\xrightarrow[\lambda\rightarrow+\infty]{}I_{p}.$ (A.6)
@@
Moreover, according to Lemma 3.2 $\kappa(\lambda)$ and $\kappa^{\prime}(\lambda)$ admit limits when $\lambda$ tends to 0. Thus we finish this proof by getting considering the definitions of $\kappa(0^{+})$ and $\kappa^{\prime}(0^{+})$ from Lemma 3.2.
@@
Consider first the function
$\hat{g}(\lambda)=\hat{r}_{\lambda}^{test}(X_{n})=\sigma^{2}+\frac{\sigma^{2}}{n}\mathrm{Tr}\big[\widetilde{S}_{p}\Delta[Q_{p}(-\lambda)]\big]+\lambda\left(\frac{\lambda\alpha^{2}}{p}-\frac{\sigma^{2}}{n}\right)\mathrm{Tr}\big[\widetilde{S}_{p}\Delta[Q_{p}^{\prime}(-\lambda)]\big].$
Since $Q_{p}(z)=(\widehat{\Sigma}_{n}-zI_{p})^{-1}$ is the resolvent of the matrix $\widehat{\Sigma}_{n}$, one has that $Q^{\prime}_{p}(z)=Q_{p}(z)^{2}$ and $Q^{\prime\prime}_{p}(z)=2Q_{p}(z)^{3}$. Consequently, the function $\hat{g}$ is differentiable for $\lambda>0$, and we obtain that
$\hat{g}^{\prime}(\lambda)=2\Big(\frac{\lambda\alpha^{2}}{p}-\frac{\sigma^{2}}{n}\Big)\mathrm{Tr}\big[\widetilde{S}_{p}\Delta[\widehat{\Sigma}_{n}Q_{p}(\lambda)^{3}]\big]$
using standard differential calculus. Since the symmetric and positive semi-definite matrices $\widehat{\Sigma}_{n}$ and $Q_{p}(\lambda)^{3}$ are commuting, it follows that $\widetilde{S}_{p}\Delta[\widehat{\Sigma}_{n}Q_{p}(\lambda)^{3}]$ is also a symmetric and positive semi-definite matrix. Thus $\mathrm{Tr}\big[\widetilde{S}_{p}\Delta[\widehat{\Sigma}_{n}Q_{p}(\lambda)^{3}]\big]\geq 0$ which means that the sign of $\hat{g}^{\prime}(\lambda)$ only depends on $\Big(\frac{\lambda\alpha^{2}}{p}-\frac{\sigma^{2}}{n}\Big)$. Indeed, if $\lambda\leq\lambda_{\ast}$ then $\hat{g}^{\prime}(\lambda)\leq 0$, and if $\lambda\geq\lambda_{\ast}$ then $\hat{g}^{\prime}(\lambda)\geq 0$. This proves that $\lambda\mapsto\hat{r}_{\lambda}^{test}(X_{n})$ attains its minimum at $\lambda=\lambda_{\ast}=\frac{\sigma^{2}p}{\alpha^{2}n}$.
Now, let us consider the function $g(\lambda)=r^{\mathrm{test}}_{\lambda}(X_{n}).$ This function is differentiable for all $\lambda>0$ and can be written as
$g(\lambda)=\sigma^{2}+\frac{\sigma^{2}}{n}\operatorname{Tr}\!\big[\widetilde{S}_{p}T_{p}(-\lambda)\big]+\lambda\Big(\frac{\lambda\alpha^{2}}{p}-\frac{\sigma^{2}}{n}\Big)\operatorname{Tr}\!\big[\widetilde{S}_{p}T_{p}^{\prime}(-\lambda)\big].$
By [HLN07][Theorem 2.4], there exists a positive matrix-valued measure $\mu=(\mu_{ij})_{1\leq i,j\leq p}\quad\text{on }\mathbb{R}^{+},$ satisfying $\mu(\mathbb{R}^{+})=I_{p}$, such that
$T_{p}(z)=\int_{0}^{+\infty}\frac{\mu(dw)}{w{\color[rgb]{0,0,0}-z}},\mbox{ for }z\in\mathbb{C}\setminus\mathbb{R}^{+}.$
Since the mapping $\color[rgb]{0,0,0}z\mapsto(w-z)^{-1}$} is differentiable for all $w\in\mathbb{R}^{+}$ and its derivatives are uniformly bounded for $\color[rgb]{0,0,0}z\in\mathbb{C}\setminus\mathbb{R}^{+}$}, differentiation under the integral sign is justified. Consequently,
$T_{p}^{\prime}(z)=\int_{0}^{+\infty}\frac{\mu(dw)}{(w-z)^{2}},\qquad T_{p}^{\prime\prime}(z)=\int_{0}^{+\infty}\frac{2\,\mu(dw)}{(w-z)^{3}},\mbox{ for }z\in\mathbb{C}\setminus\mathbb{R}^{+}.$
Differentiating $g$ yields
$g^{\prime}(\lambda)$ $=-\frac{\sigma^{2}}{n}\operatorname{Tr}\!\big[\widetilde{S}_{p}T_{p}^{\prime}(-\lambda)\big]+{\color[rgb]{0,0,0}\Big(2\frac{\lambda\alpha^{2}}{p}-\frac{\sigma^{2}}{n}\Big)}\operatorname{Tr}\!\big[\widetilde{S}_{p}T_{p}^{\prime}(-\lambda)\big]-\lambda\Big(\frac{\lambda\alpha^{2}}{p}-\frac{\sigma^{2}}{n}\Big)\operatorname{Tr}\!\big[\widetilde{S}_{p}T_{p}^{\prime\prime}(-\lambda)\big]$
$=2\Big(\frac{\lambda\alpha^{2}}{p}-\frac{\sigma^{2}}{n}\Big)\operatorname{Tr}\!\Bigg[\widetilde{S}_{p}\int_{0}^{+\infty}\Big(\frac{1}{(w+\lambda)^{2}}-\frac{\lambda}{(w+\lambda)^{3}}\Big)\,\mu(dw)\Bigg]$
$=2\Big(\frac{\lambda\alpha^{2}}{p}-\frac{\sigma^{2}}{n}\Big)\operatorname{Tr}\!\Bigg[\widetilde{S}_{p}\int_{0}^{+\infty}\frac{w}{(w+\lambda)^{3}}\,\mu(dw)\Bigg].$
Since $\mu(\mathbb{R}^{+})=I_{p}$, the matrix $\int_{0}^{+\infty}\frac{w}{(w+\lambda)^{3}}\,\mu(dw)$ is diagonal with $i$-th diagonal entry
$\int_{0}^{+\infty}\frac{w}{(w+\lambda)^{3}}\,\mu_{ii}(dw).$
As $\mu_{ii}$ is a positive measure and $\frac{w}{(w+\lambda)^{3}}>0$ for all $w>0$, each diagonal entry is strictly positive. Hence, the matrix is positive semi-definite. Since $\widetilde{S}_{p}$ is also positive semi-definite, we conclude that $\operatorname{Tr}\!\Bigg[\widetilde{S}_{p}\int_{0}^{+\infty}\frac{w}{(w+\lambda)^{3}}\,\mu(dw)\Bigg]\geq 0.$ Therefore, the sign of $g^{\prime}(\lambda)$ depends solely on the factor $\frac{\lambda\alpha^{2}}{p}-\frac{\sigma^{2}}{n},$ and we conclude as before for $\hat{g}$, that the function $g(\lambda)=r^{\mathrm{test}}_{\lambda}(X_{n})$ has a unique minimizer at $\lambda_{\ast}$. This concludes the proof of Corollary 3.1. ∎
@@ A.4 Random matrices with a variance profile in free probability
and it is sometimes defined as $\mathcal{G}_{A}$. The operator-valued $\mathcal{R}$-transform $\mathcal{R}_{A}$ of a Hermitian matrix $A$ is the unique analytic map satisfying,
@@
for all $\mathcal{Z}$ in $\mathbb{D}_{p}(\mathbb{C})^{-}$ whose diagonal entries have imaginary parts large enough in absolute value. An efficient way to produce a deterministic equivalent for a random matrix $A$ is then to find a simple approximation $\mathcal{R}_{A}^{\square}$ of $\mathcal{R}_{A}$. In this manner, one can then approximate the operator-valued Stieltjes transform of $A$ by the solution of the fixed point equation Definition A.9 where $\mathcal{R}_{A}$ is replaced by $\mathcal{R}_{A}^{\square}$. This method also allows us to compute the spectrum of perturbations of $A$ by independent matrices, see for instance [BM20].
For a generalized Wigner matrix $W_{p}$ of size $p\times p$ with a symmetric variance profile $\Gamma_{p}$, Shlyakhtenko proves [Shl96] that a good approximation of $\mathcal{R}_{W_{p}}$ is the deterministic linear map $\mathcal{R}_{W_{p}}^{\square}(\mathcal{Z})=\mathrm{deg}\left(\frac{1}{p}\Gamma_{p}\mathcal{Z}\right),\quad\mbox{for all}\quad\mathcal{Z}\in\mathbb{D}_{p}(\mathbb{C})^{-},$ where for a matrix $A$, we denote by $\mathrm{deg}(A)$ the diagonal matrix, whose $k$-diagonal element is the sum of the entries of the $k$-row of $A$.
To the best of our knowledge, apart from generalized Wigner matrices, there does not exist any other class of random matrices for which a simple approximation of the diagonal-valued $\mathcal{R}$-transform is known yet. Nevertheless, we ca now remark that, for $\mathcal{Z}=zI_{p}$ with $z\in\mathbb{C}\setminus{\mathbb{R}}^{+}$, the diagonal matrix $T_{p}(z)$ solution of the Dyson equation Equation (2.2) can be written as satisfying the fixed-point equation $T_{p}(z)=\left(-zI_{p}+\mathcal{R}_{\Sigma_{n}}^{\square}(-T_{p}(z))\right)^{-1},$ where $\mathcal{R}_{\Sigma_{n}}^{\square}$ is the non-linear map $\mathcal{R}_{\Sigma_{n}}^{\square}(\mathcal{Z})=\mathrm{deg}\left(\frac{1}{n}\Gamma_{n}^{\top}\left[I_{n}-\mathrm{deg}\left(\frac{1}{n}\Gamma_{n}\mathcal{Z}\right)\right]^{-1}\right),\quad\mbox{for all}\quad\mathcal{Z}\in\mathbb{D}_{p}(\mathbb{C})^{-}.$ This suggests that $\mathcal{R}_{\Sigma_{n}}^{\square}$ is indeed a simple approximation of the operator-valued $\mathcal{R}$-transform of $\Sigma_{n}$.
Acknowledgments
This work was supported by ANR-STARS funded project.